OpenStudy (anonymous):

find the limit. lim (x-5)/((x^2)-25) x-> 5+ i got 1/10 when i computed the limit but when i graph it, i don't see how it's 1/10...

OpenStudy (anonymous):

? should I make this for you? It's cheating

OpenStudy (anonymous):

oh! i didn't notice the point x=1/10 doesn't exist on the equation's graph ... so the empty spot that both sides of the line approaches is the limit..right?

OpenStudy (anonymous):

=lim(x-5)/[(x-5)(x+5)} =lim 1/(x+5) = 1/10